# Misiurewicz points and subhyperbolicity in unicritical algebraic correspondences

### v1

@@ 3. Background on global analytic functions
For a given $z\in\mathbb{C},$ let $S_{z}$ denote the space of all holomorphic maps locally defined in a neighborhood of $z$, with values in $\mathbb{C}.$ We define an equivalence relation on $S_{z}$ by setting $f\sim g$ provided $f$ and $g$ coincide on some small neighborhood of $z.$ Every element of $\mathfrak{S}_{z}=S_{z}/\mathord{\sim}$ is a germ of analytic function at $z.$ Let $U\subset\mathbb{C}$ be a region. Then $\mathfrak{S}(U)=\cup_{z\in U}\mathfrak{S}_{z}$ is the sheaf of germs of analytic functions on $U,$ and the elements of $\mathfrak{S}_{U}$ are denoted by $\mathfrak{f},\mathfrak{g},\ldots$ If we need to specify that $\mathfrak{f}$ is a germ at $z$ represented by some holomorphic map $f,$ then we write $\mathfrak{f}=(f,z).$ Sometimes we denote $\mathfrak{S}(U)$ by $\mathfrak{S}_{U}.$ The sheaf $\mathfrak{S}_{U}$ is a topological space: if $\mathfrak{f}=(f,z)$, then a local basis at $\mathfrak{f}$ is determined by all sets of the form $\tilde{V}=\{(f,\zeta):\zeta\in V\}$, where $V$ is a region contained in the domain of $f$ and $z\in V.$ We define the continuous map $\pi:\mathfrak{S}_{U}\to U$ by $\pi(\mathfrak{f})=z$ if $\mathfrak{f}=(f,z).$
3.2. Global analytic functions.
@@
The connected components $\mathfrak{S}_{\alpha}(U)$ of $\mathfrak{S}(U)$ are Riemann surfaces. A local chart at $\mathfrak{f}\in\mathfrak{S}_{\alpha}(U)$ is simply a restriction of the projection $\pi.$ The restriction of $\pi$ to $\mathfrak{S}_{\alpha}(U)$ is denoted by $\pi_{\alpha}.$ The map $\tilde{\mathbf{f}}_{\alpha}:\mathfrak{S}_{\alpha}(U)\to\mathbb{C}$ which assigns every germ $(f,z)$ to $f(z)$ is evidently holomorphic on the Riemann surface $\mathfrak{S}_{\alpha}(U).$ By definition, each $\tilde{\mathbf{f}}_{\alpha}$ is global analytic function, and every germ of analytic function $\mathfrak{f}$ determines a unique global analytic function $\tilde{\mathbf{f}}_{\alpha}$ such that $\mathfrak{f}\in\mathfrak{S}_{\alpha}(U).$
3.3. Analytic continuation.
@@
Given a continuous curve $\gamma:[a,b]\to U$ and a germ $\mathfrak{f}$ of holomorphic function at $\gamma(a)$, an analytic continuation of $\mathfrak{f}$ along $\gamma$ is a continuous curve $\tilde{\gamma}:[a,b]\to\mathfrak{S}(U)$ that projects onto $\gamma,$ that is, $\pi(\tilde{\gamma}(t))=\gamma(t)$, for every $t.$ It is not necessary to explicit the $\tilde{\mathbf{f}}_{\alpha}$ which is related to the continuation $\tilde{\gamma}$, for then $\{\tilde{\gamma}\}$ must be contained in a unique component $\mathfrak{S}_{\alpha}(U).$ Two analytic continuations along the same curve are either identical or differ for every $t.$ If the analytic continuation exists, then the first germ $\mathfrak{g}_{1}=\tilde{\gamma}(a)$ uniquely determines the analytic continuation along $\gamma.$ Hence $\tilde{\gamma}$ is the analytic continuation of $\mathfrak{g}_{1}$ along $\gamma.$ In the same situation, we can say that $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}=\tilde{\gamma}(b)$ are joined by an analytic continuation or that the continuation of $\mathfrak{g}_{1}$ along $\gamma$ leads to the germ $\mathfrak{g}_{2}$ at $\gamma(b).$ For example, every pair of germs in a component $\mathfrak{S}_{\alpha}(U)$ can be joined by an analytic continuation $\tilde{\gamma}$.
@@
for every $t\in[a,b]$, where $f_{t}$ is a holomorphic map locally defined at $\gamma(t)$ such that $\tilde{\gamma}(t)=(f_{t},\gamma(t))$ is an analytic continuation of $\mathfrak{f}$ along $\gamma$ with $\tilde{\gamma}(a)=\mathfrak{f}.$ Equivalently, we may also say that the analytic continuation of $\mathfrak{f}$ along $\gamma$ yields $\zeta.$
3.4. Holomorphic multifunctions.
By a holomorphic multifunction $\mathbf{f}$ we mean a set valued function that sends every point $z$ of $U$ to a subset $\mathbf{f}(z)$ of $\mathbb{C}$ in such a way that, whenever $w_{0}\in\mathbf{f}(z_{0})$, there exists a holomorphic map $f:V\to\mathbb{C}$ defined on a region, called a branch of $\mathbf{f}$, such that $f(z_{0})=w_{0}$ and $f(z)\in\mathbf{f}(z),$ for every $z\in V.$ If we want to explicit the domain of $\mathbf{f}$, we write $\mathbf{f}:U\to\mathbb{C}.$ (The usage of boldface letters avoids misinterpretations with single-valued holomorphic maps).
Any global analytic function $\tilde{\mathbf{f}}_{\alpha}$ gives rise to a holomorphic multifunction $\mathbf{f}_{\alpha}(z)=\tilde{\mathbf{f}}_{\alpha}(\pi_{\alpha}^{-1}(z))$ defined on $U.$ Given a holomorphic multifunction $\mathbf{g}:U\to\mathbb{C}$, the associated space of germs $\mathfrak{S}(\mathbf{g})$ consists of all germs $(g,z)$ such that $g$ is a branch of $\mathbf{g}$ and $z$ is in the domain of $\mathbf{g}.$ It should be noticed that every branch is holomorphic and its domain is connected.
@@
A holomorphic multifunction $\mathbf{g}:U\to\mathbb{C}$ is separable if there exists a locally finite set of exceptional points $E=E({\mathbf{g}})\subset U$ such that any point $x\in U{\setminus}B$ has a connected neighborhood $V\subset U{\setminus}E$ which is the domain of a family $\mathcal{F}$ of branches $f:V\to\mathbb{C}$, with pairwise disjoint images, such that any other local branch $h:W\to\mathbb{C}$ defined on a region $W\subset V$ is the restriction $h=f|_{W},$ for some $f\in\mathcal{F}.$
@@
It suffices to show that $\mathfrak{S}(\mathbf{f}_{\alpha})\subset\mathfrak{S}_{\alpha}(U)$. The other inclusion is immediate. We will show that every $\mathfrak{g}\in\mathfrak{S}(\mathbf{f}_{\alpha})$ belongs to $\mathfrak{S}_{\alpha}(U).$ Without loss of generality, we may assume that $\mathfrak{g}=(g,z_{0})$, for some holomorphic branch $g:\Omega\to\mathbb{C}$ of $\mathbf{f}_{\alpha}$ defined on a region $\Omega\subset U$ such that $z_{0}$ is the only point of $E$ in $\Omega$ (see Definition 3.3). Choose $z_{1}\neq z_{0}$ in $\Omega.$ Since $\mathbf{f}_{\alpha}$ is separable, there exists a connected neighborhood $V\subset\Omega{\setminus}\{z_{0}\}$ of $z_{1}$ which is the domain of a family $\mathcal{F}$ of branches of $\mathbf{f}_{\alpha}$ with pairwise disjoint images satisfying the conditions of Definition 3.3. Hence $\mathbf{f}_{\alpha}(V)$ is given by the disjoint union of all $f(V)$ such that $f\in\mathcal{F}.$ Since $g(V)$ is connected, it must be contained in some $f_{j}(V)$, with $f_{j}\in\mathcal{F}.$ This implies $g(z)=f_{j}(z)$, for every $z\in V.$ Let $S_{g}$ be the subset of the sheaf space given by all $(g,z)$ such that $z\in\Omega.$ Clearly, this set is connected in the sheaf topology, and we will show that it intersects $\mathfrak{S}_{\alpha}(U),$ which is enough to conclude that $S_{g}$ is contained in $\mathfrak{S}_{\alpha}(U).$ In order to prove this assertion, we first notice that $(f_{j},z)$ is a germ in $\mathfrak{S}_{\alpha}(U)$, for every $z\in V.$ Indeed, using the defintion of $\mathbf{f}_{\alpha}=\tilde{\mathbf{f}}_{\alpha}\circ\pi_{\alpha}^{-1}$ and the fact that $f_{j}$ is a branch of $\mathbf{f}_{\alpha}$, for any $z\in V$ there exists a holomorphic map $g_{z}$ locally defined at $z$ such that $(g_{z},z)\in\mathfrak{S}_{\alpha}(U)$ and $f_{j}(z)=g_{z}(z).$ Let $W_{z}\subset V$ be a connected neighborhood of $z$ which is contained in the domain of $g_{z}.$ Again, it follows from the definition of $\mathbf{f}_{\alpha}$ that $g_{z}:W_{z}\to\mathbb{C}$ is a branch of $\mathbf{f}_{\alpha}.$ Since $\mathbf{f}_{\alpha}$ is separable, $g_{z}|_{W_{z}}=h_{z}|_{W_{z}}$, for some $h_{z}\in\mathcal{F}$. The maps in $\mathcal{F}$ have pairwise disjoint images over $V,$ and since $h_{z}(z)=f_{j}(z),$ it follows that $h_{z}=f_{j}$, for any $z\in V.$ Hence
@@
3.5. Examples.
The elementary multifunction $\log(z)$ is a global analytic multifunction defined on $\mathbb{C}^{*}.$ Every irreducible complex polynomial in two complex variables $P(z,w)$ determines a global analytic multifunction $\mathbf{g}(z)=\{w:P(w,z)=0\}.$ In this case we have to remove from $\mathbb{C}$ a locally finite set of (algebraic) singularities (see page 3.6). Any $\mathbf{g}$ arising from some $P(z,w)$ in this way is an algebraic function. [^1] More appropriately, an algebraic multifunction. However, the term algebraic function is standard, see [Ahlfors, p. 300] If $P(z,w)$ is not irreducible then $\mathbf{g}$ is no longer a global analytic multifunction. Nevertheless, $\mathfrak{S}(\mathbf{g})$ is the union of finitely many connected components $\Lambda_{i}$ of the sheaf $\mathfrak{S}_{U}$, where $U=\mathbb{C}{\setminus}B$ and $B$ is the set of singularities. This is the case of the correspondence $(w-c)^{2}=z^{4},$ which is completely determined by two global branches: $w=z^{2}+c$ and $w=-z^{2}+c.$ The singularity at 0 is removable in this case.
@@
Suppose that $\mathbf{g}:U\to\mathbb{C}$ is a separable holomorphic multifunction with $E(\mathbf{g})=\emptyset$ and let $(\gamma,\eta)$ be a pair of continuous curves defined on $[a,b]$, with $\{\gamma\}\subset U$ and $\eta(t)\in\mathbf{g}(\gamma(t))$, for every $t\in[a,b].$ Then for every $t$ there exists a unique germ $\mathfrak{f}_{t}=(f_{t},\gamma(t))$ determined by a local branch $f_{t}$ of $\mathbf{g}$ at $\gamma(t)$ which sends $\gamma(t)$ to $\eta(t).$ Moreover, the curve $\tilde{\gamma}(t)=\mathfrak{f}_{t}$ is continuous on $[a,b]$. Hence $\eta$ is given by the analytic continuation of $\mathfrak{f}_{a}$ along $\gamma.$
@@
Since $\{\gamma\}$ is compact and $E=\emptyset$, the following property holds for some $\epsilon>0$: for any $t\in[0,1],$ there exists a unique local branch $f_{t}$ defined on the open ball $D_{t}$ centered at $\gamma(t)$, with radius $\epsilon$, such that $f_{t}$ sends $\gamma(t)$ to $\eta(t)$. There exists $\delta>0$ such that $D_{t}\cap D_{s}$ is a region and $f_{t}=f_{s}$ on $D_{t}\cap D_{s}$ if $|t-s|<\delta$ on $[0,1].$ Hence $\tilde{\gamma}(t)=(f_{t},\gamma(t))$ is a continuous path $0,1]\to\mathfrak{S}(U),$ which is evidently an analytic continuation satisfying the conditions of the statement. ∎
@@
Let $U\subset\mathbb{C}$ be a region. Assume that $\mathfrak{S}_{\alpha}(U)$ is a component of the sheaf space such that any germ $\mathfrak{f}$ in $\mathfrak{S}_{\alpha}(U)$ can be continued along any curve in $U$ starting at $\pi(\mathfrak{f}).$ Let $\gamma_{1}$ and $\gamma_{2}$ be homotopic curves in $U$, defined on $[a,b]$, with $\gamma_{1}(a)=\gamma_{2}(a)$ and $\gamma_{1}(b)=\gamma_{2}(b).$ Then the continuations of a given $\mathfrak{f}\in\mathfrak{S}_{\alpha}(U)$ with $\pi(\mathfrak{f})=\gamma_{1}(a)$ along $\gamma_{1}$ and $\gamma_{2}$ lead to the same germ at $\gamma_{1}(b).$
@@
Under the same hypothesis of Theorem 3.3, if $U$ is simply connected, then for any germ $\mathfrak{f}\in\mathfrak{S}_{\alpha}(U)$ there exists a branch $f:U\to Y$ such that $\mathfrak{f}=(f,\pi(\mathfrak{f})).$ Equivalently, every local branch can be extended to a branch defined on the whole space.
@@
Given $z\in U$, we choose some curve $\gamma$ in $U$ connecting $\pi(\mathfrak{f})$ to $z.$ The continuation of $\mathfrak{f}$ along $\gamma$ leads to a germ $\mathfrak{g}_{z}=(g_{z},z)$. By the Monodromy Theorem, $\mathfrak{g}_{z}$ does not depend on the choice of $\gamma,$ and we may define $f(z)=g_{z}(z).$ Since a small perturbation of $z$ produces essentially the same map $g_{z}$, we have a global branch $f:U\to Y$ with $\mathfrak{f}=(f,\pi(\mathfrak{f})).$
3.6. Singularities.
@@
If the continuation of every germ of $\mathbf{h}$ at $x_{j}+r$ along $\gamma$ leads back to itself, then for every germ $\mathfrak{h}$ of $\mathbf{h}$ with $\pi(\mathfrak{h})\in A_{r,j}$ there exists a unique branch $F$ of $\mathbf{h}$ defined on $A_{r,j}$ such that $F$ defines the germ $\mathfrak{h}$ at $\pi(\mathfrak{h})$, that is, $(F,\pi(\mathfrak{h}))=\mathfrak{h}.$ In other words, local branches within $A_{r,j}$ can be extended to a punctured neighborhood of the singularity. Let $\mathcal{B}_{j}$ denote the space of all branches $F$ defined on $A_{r,j}$ obtained in this way.
@@
See [Ahlfors, p. 297-300]. On page 299 of [Ahlfors] Ahlfors describes ordinary algebraic singularities. Each $x_{j}$ in Theorem 3.4 is an ordinary algebraic singularity; the function $F\in\mathcal{B}_{j}$ is obtained by analytic continuation on page 298 by $F(\zeta)=f(\zeta^{h}),$ and is the same Laurent development with $h=1$ which is present on page 299 of [Ahlfors]. ∎

### v2

@@ 3. The sheaf of germs of analytic functions
For a given $z\in\mathbb{C},$ let $S_{z}$ denote the space of all holomorphic maps locally defined in a neighborhood of $z$, with values in $\mathbb{C}.$ We define an equivalence relation on $S_{z}$ by setting $f\sim g$ provided $f$ and $g$ coincide on some small neighborhood of $z.$ Every element of $\mathfrak{S}_{z}=S_{z}/\mathord{\sim}$ is a germ of an analytic function at $z.$ Let $U\subset\mathbb{C}$ be a region. Then $\mathfrak{S}(U)=\cup_{z\in U}\mathfrak{S}_{z}$ is the sheaf of germs of analytic functions on $U,$ and the elements of $\mathfrak{S}_{U}$ are denoted by $\mathfrak{f},\mathfrak{g},\ldots$ If we need to specify that $\mathfrak{f}$ is a germ at $z$ represented by some holomorphic map $f,$ then we write $\mathfrak{f}=(f,z).$ Sometimes we denote $\mathfrak{S}(U)$ by $\mathfrak{S}_{U}.$ The sheaf $\mathfrak{S}_{U}$ is a topological space: if $\mathfrak{f}=(f,z)$, then a local basis at $\mathfrak{f}$ is determined by all sets of the form $\tilde{V}=\{(f,\zeta):\zeta\in V\}$, where $V$ is a region contained in the domain of $f$ and $z\in V.$ We define the continuous map $\pi:\mathfrak{S}_{U}\to U$ by $\pi(\mathfrak{f})=z$ if $\mathfrak{f}=(f,z).$
Global analytic functions.
@@
The connected components $\mathfrak{S}_{\alpha}(U)$ of $\mathfrak{S}(U)$ are Riemann surfaces. A local chart at $\mathfrak{f}\in\mathfrak{S}_{\alpha}(U)$ is simply a restriction of the projection $\pi.$ The restriction of $\pi$ to $\mathfrak{S}_{\alpha}(U)$ is denoted by $\pi_{\alpha}.$ The map $\tilde{\mathbf{f}}_{\alpha}:\mathfrak{S}_{\alpha}(U)\to\mathbb{C}$ which assigns every germ $(f,z)$ to $f(z)$ is evidently holomorphic on the Riemann surface $\mathfrak{S}_{\alpha}(U).$ By definition, each $\tilde{\mathbf{f}}_{\alpha}$ is a global analytic function, and every germ of analytic function $\mathfrak{f}$ determines a unique global analytic function $\tilde{\mathbf{f}}_{\alpha}$ such that $\mathfrak{f}\in\mathfrak{S}_{\alpha}(U).$
Analytic continuation.
@@
Given a continuous curve $\gamma:[a,b]\to U$ and a germ $\mathfrak{f}$ of a holomorphic function at $\gamma(a)$, an analytic continuation of $\mathfrak{f}$ along $\gamma$ is a continuous curve $\tilde{\gamma}:[a,b]\to\mathfrak{S}(U)$ that projects onto $\gamma,$ that is, $\pi(\tilde{\gamma}(t))=\gamma(t)$, for every $t.$ It is not necessary to explicitly describe the $\tilde{\mathbf{f}}_{\alpha}$ associated with the continuation $\tilde{\gamma}$, since $\{\tilde{\gamma}\}$ must be contained in a unique component $\mathfrak{S}_{\alpha}(U).$ Two analytic continuations along the same curve are either identical or differ for every $t.$ If the analytic continuation exists, then the first germ $\mathfrak{g}_{1}=\tilde{\gamma}(a)$ uniquely determines the analytic continuation along $\gamma.$ Hence $\tilde{\gamma}$ is the analytic continuation of $\mathfrak{g}_{1}$ along $\gamma.$ In the same situation, we can say that $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}=\tilde{\gamma}(b)$ are joined by an analytic continuation or that the continuation of $\mathfrak{g}_{1}$ along $\gamma$ leads to the germ $\mathfrak{g}_{2}$ at $\gamma(b).$ For example, every pair of germs in a component $\mathfrak{S}_{\alpha}(U)$ can be joined by an analytic continuation $\tilde{\gamma}$.
@@
for every $t\in[a,b]$, where $f_{t}$ is a holomorphic map locally defined at $\gamma(t)$ such that
$\tilde{\gamma}(t)=(f_{t},\gamma(t))$
is an analytic continuation of $\mathfrak{f}$ along $\gamma$ with $\tilde{\gamma}(a)=\mathfrak{f}.$ Equivalently, we may also say that the analytic continuation of $\mathfrak{f}$ along $\gamma$ yields $\zeta.$
Holomorphic multifunctions.
By a holomorphic multifunction $\mathbf{f}$ we mean a set-valued function that sends every point $z$ of $U$ to a subset $\mathbf{f}(z)$ of $\mathbb{C}$ in such a way that, whenever $w_{0}\in\mathbf{f}(z_{0})$, there exists a holomorphic map $f:V\to\mathbb{C}$ defined on a region, called a branch of $\mathbf{f}$, such that $f(z_{0})=w_{0}$ and $f(z)\in\mathbf{f}(z),$ for every $z\in V.$ If we wish to make the domain of $\mathbf{f}$ explicit, we write $\mathbf{f}\colon U\to\mathbb{C}.$
(The usage of boldface letters avoids misinterpretations with single-valued holomorphic maps).
Any global analytic function $\tilde{\mathbf{f}}_{\alpha}$ gives rise to a holomorphic multifunction
$\mathbf{f}_{\alpha}(z)=\tilde{\mathbf{f}}_{\alpha}(\pi_{\alpha}^{-1}(z))$
defined on $U.$ Given a holomorphic multifunction $\mathbf{g}:U\to\mathbb{C}$, the associated space of germs $\mathfrak{S}(\mathbf{g})$ consists of all germs $(g,z)$ such that $g$ is a branch of $\mathbf{g}$ and $z$ is in the domain of $\mathbf{g}.$ It should be noticed that every branch is holomorphic and its domain is connected.
@@
A holomorphic multifunction $\mathbf{g}:U\to\mathbb{C}$ is separable if there exists a locally finite set of exceptional points $E=E({\mathbf{g}})\subset U$ such that any point $x\in U{\setminus}E$ has a connected neighborhood $V\subset U{\setminus}E$ which is the domain of a family $\mathcal{F}$ of branches $f:V\to\mathbb{C}$, with pairwise disjoint images, such that any other local branch $h:W\to\mathbb{C}$ defined on a region $W\subset V$ is the restriction $h=f|_{W},$ for some $f\in\mathcal{F}.$
@@
It suffices to show that $\mathfrak{S}(\mathbf{f}_{\alpha})\subset\mathfrak{S}_{\alpha}(U)$. The other inclusion is immediate. We will show that every $\mathfrak{g}\in\mathfrak{S}(\mathbf{f}_{\alpha})$ belongs to $\mathfrak{S}_{\alpha}(U).$ Without loss of generality, we may assume that $\mathfrak{g}=(g,z_{0})$, for some holomorphic branch $g:\Omega\to\mathbb{C}$ of $\mathbf{f}_{\alpha}$ defined on a region $\Omega\subset U$ such that $z_{0}$ is the only point of $E$ in $\Omega$ (see Definition 3.3). Choose $z_{1}\neq z_{0}$ in $\Omega.$ Since $\mathbf{f}_{\alpha}$ is separable, there exists a connected neighborhood $V\subset\Omega{\setminus}\{z_{0}\}$ of $z_{1}$ which is the domain of a family $\mathcal{F}$ of branches of $\mathbf{f}_{\alpha}$ with pairwise disjoint images satisfying the conditions of Definition 3.3. Hence $\mathbf{f}_{\alpha}(V)$ is given by the disjoint union of all $f(V)$ such that $f\in\mathcal{F}.$ Since $g(V)$ is connected, it must be contained in some $f_{j}(V)$, with $f_{j}\in\mathcal{F}.$ This implies $g(z)=f_{j}(z)$, for every $z\in V.$ Let $S_{g}$ be the subset of the sheaf given by all $(g,z)$ such that $z\in\Omega.$ Clearly, this set is connected in the sheaf topology, and we will show that it intersects $\mathfrak{S}_{\alpha}(U),$ which is enough to conclude that $S_{g}$ is contained in $\mathfrak{S}_{\alpha}(U).$ In order to prove this assertion, we first notice that $(f_{j},z)$ is a germ in $\mathfrak{S}_{\alpha}(U)$, for every $z\in V.$ Indeed, using the definition of $\mathbf{f}_{\alpha}=\tilde{\mathbf{f}}_{\alpha}\circ\pi_{\alpha}^{-1}$ and the fact that $f_{j}$ is a branch of $\mathbf{f}_{\alpha}$, for any $z\in V$ there exists a holomorphic map $g_{z}$ locally defined at $z$ such that $(g_{z},z)\in\mathfrak{S}_{\alpha}(U)$ and $f_{j}(z)=g_{z}(z).$ Let $W_{z}\subset V$ be a connected neighborhood of $z$ which is contained in the domain of $g_{z}.$ Again, it follows from the definition of $\mathbf{f}_{\alpha}$ that $g_{z}:W_{z}\to\mathbb{C}$ is a branch of $\mathbf{f}_{\alpha}.$ Since $\mathbf{f}_{\alpha}$ is separable, $g_{z}|_{W_{z}}=h_{z}|_{W_{z}}$, for some $h_{z}\in\mathcal{F}$. The maps in $\mathcal{F}$ have pairwise disjoint images over $V,$ and since $h_{z}(z)=f_{j}(z),$ it follows that $h_{z}=f_{j}$, for any $z\in V.$ Hence
@@
Examples.
The elementary multifunction $\log(z)$ is a global analytic multifunction defined on $\mathbb{C}^{*}.$ Every irreducible complex polynomial in two complex variables $P(z,w)$ determines a global analytic multifunction $\mathbf{g}(z)=\{w:P(w,z)=0\}.$ In this case we have to remove from $\mathbb{C}$ a locally finite set of (algebraic) singularities (see page 3). Any $\mathbf{g}$ arising from some $P(z,w)$ in this way is an algebraic function. [^1] More appropriately, an algebraic multifunction. However, the term algebraic function is standard, see [1, p. 300] If $P(z,w)$ is not irreducible then $\mathbf{g}$ is no longer a global analytic multifunction. Nevertheless, $\mathfrak{S}(\mathbf{g})$ is the union of finitely many connected components $\Lambda_{i}$ of the sheaf $\mathfrak{S}_{U}$, where $U=\mathbb{C}{\setminus}B$ and $B$ is the set of singularities. This is the case of the correspondence $(w-c)^{2}=z^{4},$ which is completely determined by two global branches: $w=z^{2}+c$ and $w=-z^{2}+c.$ The singularity at 0 is removable in this case.
@@
Suppose that $\mathbf{g}:U\to\mathbb{C}$ is a separable holomorphic multifunction with $E(\mathbf{g})=\emptyset$ and let $(\gamma,\eta)$ be a pair of continuous curves defined on $[a,b]$, with $\{\gamma\}\subset U$ and $\eta(t)\in\mathbf{g}(\gamma(t))$, for every $t\in[a,b].$ Then for every $t$ there exists a unique germ
$\mathfrak{f}_{t}=(f_{t},\gamma(t))$
determined by a local branch $f_{t}$ of $\mathbf{g}$ at $\gamma(t)$ which sends $\gamma(t)$ to $\eta(t).$ Moreover, the curve $\tilde{\gamma}(t)=\mathfrak{f}_{t}$ is continuous on $[a,b]$. Hence $\eta$ is given by the analytic continuation of $\mathfrak{f}_{a}$ along $\gamma.$
@@
Since $\{\gamma\}$ is compact and $E=\emptyset$, the following property holds for some $\epsilon>0$: for any $t\in[a,b],$ there exists a unique local branch $f_{t}$ defined on the open ball $D_{t}$ centered at $\gamma(t)$, with radius $\epsilon$, such that $f_{t}$ sends $\gamma(t)$ to $\eta(t)$. There exists $\delta>0$ such that $D_{t}\cap D_{s}$ is a region and $f_{t}=f_{s}$ on $D_{t}\cap D_{s}$ if $|t-s|<\delta$ on $[a,b].$ Hence $\tilde{\gamma}(t)=(f_{t},\gamma(t))$ is a continuous path $a,b]\to\mathfrak{S}(U),$ which is evidently an analytic continuation satisfying the conditions of the statement. ∎
@@
Let $U\subset\mathbb{C}$ be a region. Assume that $\mathfrak{S}_{\alpha}(U)$ is a component of the sheaf such that any germ $\mathfrak{f}$ in $\mathfrak{S}_{\alpha}(U)$ can be continued along any curve in $U$ starting at $\pi(\mathfrak{f}).$ Let $\gamma_{1}$ and $\gamma_{2}$ be homotopic curves in $U$, defined on $[a,b]$, with $\gamma_{1}(a)=\gamma_{2}(a)$ and $\gamma_{1}(b)=\gamma_{2}(b).$ Then the continuations of a given $\mathfrak{f}\in\mathfrak{S}_{\alpha}(U)$ with $\pi(\mathfrak{f})=\gamma_{1}(a)$ along $\gamma_{1}$ and $\gamma_{2}$ lead to the same germ at $\gamma_{1}(b).$
@@
Under the same hypothesis of Theorem 3.3, if $U$ is simply connected, then for any germ $\mathfrak{f}\in\mathfrak{S}_{\alpha}(U)$ there exists a branch $f:U\to\mathbb{C}$ such that $\mathfrak{f}=(f,\pi(\mathfrak{f})).$ Equivalently, every local branch can be extended to a branch defined on the whole space.
@@
Given $z\in U$, we choose some curve $\gamma$ in $U$ connecting $\pi(\mathfrak{f})$ to $z.$ The continuation of $\mathfrak{f}$ along $\gamma$ leads to a germ $\mathfrak{g}_{z}=(g_{z},z)$. By the Monodromy Theorem, $\mathfrak{g}_{z}$ does not depend on the choice of $\gamma,$ and we may define $f(z)=g_{z}(z).$ Since a small perturbation of $z$ produces essentially the same map $g_{z}$, we have a global branch $f:U\to\mathbb{C}$ with $\mathfrak{f}=(f,\pi(\mathfrak{f})).$
Singularities.
@@
If the continuation of every germ of $\mathbf{h}$ at $x_{j}+r$ along $\gamma$ leads back to itself, then for every germ $\mathfrak{h}\in\mathfrak{S}(\mathbf{h})$ with $\pi(\mathfrak{h})\in A_{r,j}$ there exists a unique branch $F$ of $\mathbf{h}$ defined on $A_{r,j}$ such that $F$ defines the germ $\mathfrak{h}$ at $\pi(\mathfrak{h})$, that is, $(F,\pi(\mathfrak{h}))=\mathfrak{h}.$ In other words, local branches within $A_{r,j}$ can be extended to a punctured neighborhood of the singularity. Let $\mathcal{B}_{j}$ denote the space of all branches $F$ defined on $A_{r,j}$ obtained in this way.
@@
See [1, p. 297-300]. On page 299 of [1] Ahlfors describes ordinary algebraic singularities. Each $x_{j}$ in Theorem 3.4 is an ordinary algebraic singularity; the function $F\in\mathcal{B}_{j}$ is obtained by analytic continuation on page 298 by $F(\zeta)=f(\zeta^{h}),$ and is the same Laurent development with $h=1$ which is present on page 299 of [1]. ∎